Leslie Sporting Goods is a locally owned store that specializes in printing team jerseys. The majority of its business comes from orders for various local teams and organizations. While Leslie’s prints everything from bowling team jerseys to fraternity/sorority apparel to special event shirts, summer league baseball and softball team jerseys are the company’s biggest source of revenue.

A portion of Leslie’s operating information for the company’s last year follows:

Number of Operating
Month Jerseys Printed Cost
January 215 $5,500
February 210 5,740
March 380 5,800
April 625 8,675
May 750 9,000
June 630 9,760
July 400 6,200
August 350 6,155
September 300 5,980
October 330 6,010
November 200 4,950
December 150 4,500


Required:
3.
Using the high-low method, calculate the store’s total fixed operating costs and variable operating cost per jersey.
4. Using the high-low method results, calculate the store’s expected operating cost if it printed 480 jerseys.
5. Perform a least-squares regression analysis on Leslie’s data.
6. Using the regression output, create a linear equation (y = a + bx) for estimating Leslie’s operating costs.
7. Using the least-squares regression results, calculate the store’s expected operating cost if it prints 625 jerseys.
Answer #1

3.
Variable Cost per unit = ($9000-4500)/(750-150) = $7.50 per Jersey
Fixed Costs = $9000 - 750 x 7.5 = $3375

4.
Cost for 480 Jersey's = $3375 + 480 x $7.5 = $6975

5.

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.94925
R Square 0.901076
Adjusted R Square 0.891184
Standard Error 551.9093
Observations 12
ANOVA
df SS MS F Significance F
Regression 1 27745737 27745737 91.08794 2.43E-06
Residual 10 3046038 304603.8
Total 11 30791775
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 3414.056 362.5764 9.416103 2.75E-06 2606.186 4221.927 2606.186 4221.927
Number of Jerseys 8.216151 0.860871 9.544 2.43E-06 6.298011 10.13429 6.298011 10.13429

6. Total Cost = $3414.06 + $8.22 x Number of Jersey's

7. Cost for 625 Jerseys = $3414.06 + $8.22 x 625 = $8551.56